OpenStudy (anonymous):

Find the points on the curve y = 2x^3 - 12 x^2 + 18 x + (-3) at which the tangent line is horizontal. The points that have a horizontal tangent line are . (You are to give the points, not only the x-values. If there are more than one point, separate the points with a comma in the answer space.)

OpenStudy (anonymous):

(1,0),(3,0) thats what i got for the answer but this is incorrect

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

y=2x^3 - 12x^2 + 18x - 3 dy/dx= 6x^2 -24x +18 tangent line is horizontal, dy/dx=0 6x^2 -24x +18=0 (x-3)(x-1)=0 x=1,3 when x=1, y= 2(1^3)-12(1^2)+18(1)-3= 5 when x=3, y=-3 (1,5) and (3,-3) This is correct.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

[y=2x^3−12x^2+18x+(−3)\]\[y′=6x^2−24x+18=6(x^2−4x+3)=6(x−1)(x−3)=0→x=1,x=3\]\

OpenStudy (anonymous):

y(1)=5 y(3)=-3

OpenStudy (anonymous):

Same as above.